OpenStudy (anonymous):

Which expression completes this equation? 36 ÷ 4 – 2 • 3 = ? A. [18 – (10 – 6)] + 1 B. [(8 – 6) • 3] ÷ 2 C. 36 – 2 • 3 D. 3 • 6 – 4 ÷ 2

OpenStudy (anonymous):

[(8 - 6) * 3] / 2 = [2 * 3] / 2 = 6/2 = 3

OpenStudy (anonymous):

so B c:

OpenStudy (anonymous):

thanks. :)